# CheXalign: Preference fine-tuning in chest X-ray interpretation models without human feedback

### v2

@@ 7 Limitations
Our work focuses on only two VLMs, CheXagent (Chen et al., 2024a) and CheXagent-2 (Chen et al., 2024b). Other VLMs, from different families and sizes, may behave differently. In addition, our study of CheXagent-2 was more limited in nature and should be considered more preliminary.
Moreover, we treat clinically relevant metrics such as GREEN and CheXbert scores as the silver standard. While these metrics are highly relevant, a thorough evaluation involving clinical experts, radiologists, should be conducted for this study to be considered complete. This is an important direction for future work.
@@
Finally, we restrict ourselves to only offline DAAs. This leaves out a range of very competitive alignment algorithms, including on-policy RL algorithms, as well as the online, or iterative, counterparts to the DAAs considered. The recent success of DeepSeek-R1 DeepSeek-AI et al. (2025) has led to a resurgence of interest in on-policy RL alignment algorithms. We believe that this is an very interesting avenue for future work.
@@ B.3 Sampling Details
In this paper, we set up CheXagent to treat the cases where the two CXRs are a frontal and a lateral image (i.e. from the same point in time) and a frontal and a prior frontal image (i.e. from two points in time) as separate cases with a separate prompt. This is in contrast to CheXagent-2, which employs the same prompt for both cases. Moreover, we employed stochastic sampling with temperature $1.0$ for CheXagent at test time, while CheXagent-2 is based on greedy sampling, both with beam size set to $1$. For preference pairs generation, both employ stochastic sampling with temperature $1.0$ and beam size set to $1$ These differences will explain parts of the performance differences between CheXagent and CheXagent-2. However, this is of minor concern to this study as we are mainly interested in the effect of preference fine-tuning, comparing the same setup prior and post alignment.
@@ B.4 Evaluation Details
GREEN (Ostmeier et al., 2024) and GREEN-LC are based on the official implementation [^11] https://github.com/Stanford-AIMI/GREEN using StanfordAIMI/GREEN-radllama2-7b. The BERTScore (Zhang et al., 2019) used is from evaluate (v0.4.0) with distilbert-base-uncased as BERT model. The F1 CheXbert (Smit et al., 2020) scores use f1chexbert (v0.0.2) with default configurations. ROUGE (Lin, 2004) is from rouge-score (v0.1.2) using rougeL. Finally, the BLEU (Papineni et al., 2002) is BLEU-4, based on a custom code, available in the code base corresponding to this project.
@@ Appendix C Additional results
C.1 Summary Statistics of Preference Data

### v3

@@ 5.4 Alignment Tax Analysis
5.5 Qualitative Analysis
To further shed light on policy behavior pre and post alignment, a qualitative study was conducted by selecting three interesting cases from the MIMIC-CXR test set, covering a range of positive and negative findings and the presence and absence of indications. A board-certified radiologist was then asked to color-code candidate reports from the SFT baselines and aligned policies. Results for CheXagent-2 using GREEN as Judge for the first example are available in Fig. 4. For this case, despite being very strong, CheXagent-2 exhibited two errors: 1) it incorrectly reported clear lungs and 2) it failed to detect small bilateral pleural effusions. Notably, the aligned version addressed both of these errors. A similar conclusion is true for the other two cases. In addition, GREEN score as Judge seems to be superior to BERTScore and we observe exact, and semantically equivalent, repetitions in some cases. Full results, along with an extensive discussion, are available in §C.4.
Figure 4: Color-coded candidate reports. Green and red represent correct and incorrect. Orange highlights references to a prior imaging study.
@@ 7 Limitations
Our work focuses on only two VLMs, CheXagent (Chen et al., 2024a) and CheXagent-2 (Chen et al., 2024b). While different in terms of underlying architecture, information used, and baseline performance, it would be of interest to additionally study VLMs from other families and sizes. In addition, our study of CheXagent-2 was more limited in nature and should be considered more preliminary.
Moreover, we treat clinically relevant metrics such as GREEN and CheXbert scores as the silver standard. While these metrics are highly relevant, a thorough evaluation involving clinical experts, radiologists, should be conducted for this study to be considered complete. Our qualitative analysis is a first step in this direction. However, conducting larger-scale reader studies remains an important direction for future work. We acknowledge potential biases beyond verbosity, such as societal biases related race, sex, or other demographic factors that may be embedded either in the underlying data or in the Judge. These potential biases warrant thorough investigation and mitigation in future work.
@@
Finally, we restrict ourselves to only offline DAAs. This leaves out a range of very competitive alignment algorithms, including on-policy RL algorithms, as well as the online, or iterative, counterparts to the DAAs considered. The recent success of DeepSeek-R1 (DeepSeek-AI et al., 2025) has led to a resurgence of interest in on-policy RL alignment algorithms. In particular, developing “verification functions” for the RRG task and employing reinforcement learning with verifiable rewards (RLVR) (Lambert et al., 2025) is an very interesting avenue for future work.
8 Acknowledgements
We would like to acknowledge compute support from Microsoft, Google, and Stanford Marlowe (Kapfer et al., 2025). D.H. is supported by MedTechLabs, the Göran Gustafsson foundation, and the foundation for Technical Scientific Research and Education. J.X. acknowledges support from the Canadian Institutes of Health Research (CIHR) and Oxford University Press. S.O. acknowledges funding from the DFG (German Research Foundation, ID: 517316550). M.V. is supported by graduate fellowship awards from the Department of Defense (NDSEG), the Knight-Hennessy Scholars program at Stanford University, and the Quad program. C.B. receives research support from the Promedica Foundation, Chur, CH. A.C. receives research support from NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060; ARPA-H grants AY2AX000045 and 1AYSAX0000024-01; and NIH contracts 75N92020C00008 and 75N92020C00021. Z.C. receives support from the AIMI-AWS Cloud Credit Program and the HAI-Google research grant.
@@ B.3 Sampling Details
In this paper, we set up CheXagent to treat the cases where the two CXRs are a frontal and a lateral image (i.e. from the same point in time) and a frontal and a prior frontal image (i.e. from two points in time) as separate cases with a separate prompt. This is in contrast to CheXagent-2, which employs the same prompt for both cases. Moreover, we employed stochastic sampling with temperature 1.01.01.0 for CheXagent at test time, while CheXagent-2 is based on greedy sampling, both with beam size set to 111. For preference pairs generation, both employ stochastic sampling with temperature 1.01.01.0 and beam size set to 1.1.1 . These differences will explain parts of the performance differences between CheXagent and CheXagent-2. However, this is of minor concern to this study as we are mainly interested in the effect of preference fine-tuning, comparing the same setup prior and post alignment.
@@ B.4 Evaluation Details
GREEN (Ostmeier et al., 2024) and GREEN-LC are based on the official implementation [^12] https://github.com/Stanford-AIMI/GREEN. using StanfordAIMI/GREEN-radllama2-7b. The BERTScore (Zhang et al., 2019) used is from evaluate (v0.4.0) with distilbert-base-uncased as BERT model. The F1 CheXbert (Smit et al., 2020) scores use f1chexbert (v0.0.2) with default configurations. ROUGE (Lin, 2004) is from rouge-score (v0.1.2) using rougeL. Finally, the BLEU (Papineni et al., 2002) is BLEU-4, based on a custom code, available in the code base corresponding to this project.
@@ Appendix C Additional results
C.1 Preference data
Summary statistics for the chosen and rejected subsets are available in Table 8. We also report summary statistics of the length (in words) of the generated reports. Notably, the difference, or spread, in average length of the chosen and rejected subsets is slightly more pronounced for GREEN than for the BERTScore, 6.9% compared to 5.8% and much more significant overall for CheXagent-2 with a difference of 19.3% and 17.1%, respectively. To build further build intuition on the resulting chosen and rejected subsets we report CheXbert scores in Table 9. Scores for CheXagent and CheXagent-2 are not directly comparable since the training data is different. As expected, these scores indicate that the chosen subset is considerably better than the rejected. The average CheXbert scores for the chosen subset are more or less the same for GREEN and BERTScore as Judge. Similarly is true for the rejected subset, though now BERTScore as Judge results in slightly lower scores–meaning that the spread is marginally larger. Finally, as a qualitative study, we include a couple of examples of rejected and chosen candidates for CheXagent and CheXagent-2 is Fig. 5 and Fig. 6, respectively. We include the score assigned by the respective Judge in brackets.
@@ C.1 Preference data
Chosen Rejected
F1-14 F1-5 F1-14 F1-5
Model Judge Macro ()($\uparrow$)( ↑ ) Micro ()($\uparrow$)( ↑ ) Macro ()($\uparrow$)( ↑ ) Micro ()($\uparrow$)( ↑ ) Avg. ()($\uparrow$)( ↑ ) Macro ()($\uparrow$)( ↑ ) Micro ()($\uparrow$)( ↑ ) Macro ()($\uparrow$)( ↑ ) Micro ()($\uparrow$)( ↑ ) Avg.  ()($\uparrow$)( ↑ )
CheXagent GREEN 0.574 0.668 0.649 0.700 0.648 0.411 0.506 0.455 0.507 0.463
BERTScore 0.582 0.680 0.647 0.700 0.652 0.395 0.488 0.444 0.493 0.419
CheXagent-2 GREEN 0.488 0.587 0.550 0.621 0.561 0.336 0.416 0.365 0.415 0.373
BERTScore 0.488 0.594 0.541 0.615 0.560 0.316 0.394 0.346 0.393 0.362
Table 9: CheXbert scores on the MIMIC-CXR on the chosen and rejected subsets.
Figure 5: First two examples of chosen and rejected candidates for CheXagent in the MIMIC-CXR train set. Number in brackets is assigned score (GREEN or BERTScore).
Figure 6: First two examples of chosen and rejected candidates for CheXagent-2 in the MIMIC-CXR train set. Number in brackets is assigned score (GREEN or BERTScore).
@@ C.3 Generalization to CheXbert Scores
C.4 Qualitative Analysis
To build some further intuition of behavior pre and post alignment, we conduct a qualitative study in which a board-certified radiologist was asked to color-code candidate reports from CheXagent, CheXagent-2, as well as their top performing aligned versions for GREEN and BERTScore as Judge, respectively, in terms of cheXbert scores in Table 4. We selected three interesting, and representative, examples from MIMIC-CXR test set, including examples with positive and negative findings, only negative findings, and with and without indications. These were subsequently given to a radiologist to color-code, using the reference reports as well as the CXRs in png format (i.e. exact measurements are not possible). Three cases with six reports each, resulted in a total of 18 reports to be processed. The reports are color-coded as green (correct), red (incorrect), and orange to indicate where an reference to prior report is made though none was presented at the time of inference. For instance, the statement that something is “stable” may or may not be true, depending on the prior imaging study.
Example 1 and 2 are available in Fig. 7. For example 1, CheXagent-2, despite being a very strong model, contains two errors. First, it incorrectly states that lungs are elsewhere clear—marked in red. Secondly, comparing with the reference report, the detection of small bilateral pleural effusions is missing. None of these errors are made by the aligned versions, using both GREEN and BERTScore as Judge. Moreover, for CheXagent with both GREEN and BERTScore as Judge, there are repetitions. For GREEN “no overt edema” is mentioned twice. For BERTScore as Judge, “no pneumothorax” is repeated. These types of exact repetitions may be due to verbosity bias.
Moving on to example 2, we can see that there are exact repetition for CheXagent-2 aligned with DPO. “There is no pneumothorax” and “There is no pleural effusion” are bot repeated as “There is no evidence of …”. Interestingly, such repetition are not present for BERTScore as Judge, nor for the results from CheXagent. Notably, however, both for CheXagent-2 and CheXagent there are errors in the candidate report using BERTScore as Judge for this case. “Support devices have been removed” from CheXagent cannot cannot be supported nor refuted definitively.
Finally, consider example 3 in Fig. 8. This case includes a lot of details, with a reference report significantly longer than the previous two cases. One thing that is very interesting about this case is that the reference report includes a quantitative measurement: the distance between where the endotracheal tube terminates and the carina, expressed in centimeters. Quantitative results are given in some of the candidate reports. However, due to the set up of this reader study (i.e. that we used CXRs in png format) it is not feasible for the radiologist to make measurements and determine distance definitively. To capture this additional feature, we introduce the additional color blue. Takeaways are similar as to the previous two cases. Errors are made by CheXagent-2 and CheXagent, as well as the aligned versions using BERTScore as Judge. For GREEN as Judge, the performance remains strong, with no direct errors. Moreover, cheXagent makes a partial mistake not capture by our color coding scheme in this case as it states that there is “A small right pleural effusion,” where it is actually bilateral.
Figure 7: Color-coded candidate reports: Examples 1 and 2. Green and red represent correct and incorrect. Orange refers to prior imaging study.
Figure 8: Color-coded candidate reports: Example 3. Green and red represent correct and incorrect. Orange refers to prior imaging study. Blue indicates measurements that have not been verified.